OpenStudy (anonymous):

X^2 + 6x +8=0 solve by factoring

OpenStudy (gorv):

x^@+4x+2x+8=0 x(x+4)+2(x+4)=0 (x+4)(x+2)=0

OpenStudy (gorv):

@msbryant8